OpenStudy (anonymous):

5 + 6 x 8 = 5 + 48 I understand that 6 x 8 = 48 + 5 = 53 But you have to change them and it becomes 20 x 8 _ 4 = 160 - 4 = 156 How did it become that please explain

OpenStudy (rajee_sam):

what is your question, not clear at all??

OpenStudy (anonymous):

5 + 6 x 8 = 5 +48 = 53

OpenStudy (anonymous):

I know how to get that but then you change it and it becomes

OpenStudy (rajee_sam):

change what?

OpenStudy (anonymous):

20 x 8 - 4 =160 - 4 = 156

OpenStudy (anonymous):

Change the sum:L

OpenStudy (rajee_sam):

you know order of operations??

OpenStudy (rajee_sam):

PEMDAS

OpenStudy (anonymous):

+ becomes x and x becomes -

OpenStudy (rajee_sam):

no, if you have lot of numbers with different operations ( x, /, +, -) in between them, which one of the operation to do first??

OpenStudy (rajee_sam):

say for example I have\[5 + (9\div3) \times 15 - 12\]

OpenStudy (rajee_sam):

I have all +, /, x and - in this expression right?